Consider the following aggregate expenditure model of the Canadian economy operating with given wages and other factor prices, price level, interest rates, exchange rates, and expectations:
C = 50 + 0.8YD I = 400 G = 500 T = 0.3Y X = 650 IM = 0.36Y
where C is consumption (the 0.8 term represents the marginal propensity to consume) YD is disposable income, I is investment, G is government spending on goods and services, T is the total value of taxes net of transfers (the 0.3 term represents the net tax rate on national income), X is exports, and IM is imports (the 0.36 term represents the marginal propensity to import).
(d) Nowsupposethatthegovernmentdecidestouseitsspendingpowertorestorenationalincometoits original level. By how much must the government increase G to restore the original level of national income? What will happen to the government’s budget balance? How do you explain the new level of the budget balance compared to that in part (c) and in part (b)?
1
Expert's answer
2015-02-23T08:39:49-0500
C = 50 + 0.8YD I = 400 G = 500 T = 0.3Y X = 650 IM = 0.36Y
(a) AE = Y = C + I + G + NX = 50 + 0.8(Y - 0.3Y) + 400 + 500 + 650 - 0.36Y = 0.2Y + 1600
Y = 0.2Y + 1600
Y = 1600/0.8 = 2000
So, the equilibrium level of national income is 2000.
The value of the multiplier is:
m = 1/(1 - c*(1 - t) + im) = 1/(1 - 0.8*(1 - 0.3) + 0.36) = 1/0.8 = 1.25